Prev Next

SW&IT Meeting Minutes 2022-04-25

Attendees: Brandon, Glenn, David, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220418
LR20220418 - Bec, Matthieu 32.0
LR20220418 - Mechtley, Brandon 80.0
LR20220418 - Eychaner, Glenn 40.0
LR20220418 - Edmeades, David 40.0
LR20220418 - Grenz, Paul 18.0
LR20220418 - Hooper, Stephen 0.0
LR20220418 - Zhang, Yang 40.0
LR20220418 - Zhang, Xianyu 40.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • What distribution should we standardize on with CentOS 7 only 2 years from EOL? Some possibilities.

Team reports

MB

  • Last week
    • OC, Management, ATT, AOSWSYS
    • DMS dimm pre-2015, satellite avoidance charts
    • OT/PIT/web obsdev meeting
    • it8573-luci-hangs-after-ao-loops-closed (slack) logs investigation
  • This week
    • board, GMT M1 review
    • LBTI mongo ingest
    • mongo archive cron
    • ARGOS github synch (Lorenzo?)
    • aorev_git dx asm, others ctd.

DE

  • Monday: Jury duty, result: not selected
  • Working with SE on LBTI Google Workspace migration
    • Nothing as of this morning, I'll poke him so we don't get behind the curve on this.
  • IssueTrak notificaiton problem
    • Opened ticket with IT; they are looking into it
    • SA reported a good notification over the weekend
    • IT responded, found a bug and reported it to the dev team.
      • They also note that HTML email would avoid the problem but understand that we choose plain text on purpose
  • Mac Mini ready to go, will install ASAP

YZ

  • Previous:
    • Some discussions on TMS calculations with Andrew.
    • Working on a TMS software update (to be released soon).
    • Some issuetraks (mods1data freeze, LBC guiding, TCS shared memory).
  • Next:
    • Complete TMS software update and work with Heejoo to schedule a test.
    • Work on revise some of the TCS patching/release document.
    • Continue some of the issuetraks work (perhaps meeting with Olga for some LBC/MODS discussions).

PSG

  • Last week:
    • Look at different versions of AZCAM software
    • Talk to Gustavo about different people who should know about new ARGOS interface.
    • NOMIC, LMIRCAM testing with updated software.
    • Started updating SHARKNIR software with new buffer length parameter.
  • This week:
    • Finish SHARKNIR update.
    • Finish AZCAM software inspection.
    • Address some concerns Gustavo has about powering ON/OFF the white light source.

GE

  • Last week:
    • Finished upgrading my TCS VM to the latest version. Improved the TCS build system, added my TCS build notes to README.md, and submitted a pull request.
      • The 2021A TCS and the LUCI software can be built and run on the latest CentOS 7 release (7.9.2009). Perhaps this summer shutdown we should schedule IT time to update?
    • Discussed with Instruments and SciOps when to clean the LUCI database disk space (IT 8454), as the disk is now 90% full. Possibly May 14-21.
    • Further discussed and investigated IT 8573 (RunAO timing out), in an attempt to reach a conclusion.
    • Investigated the LUCI detector error reported 17-Apr-2022 3:49AM by B.Rothberg (IT 8603), and sent the system journal logs to R.Mather.
    • Resolved LUCI ITs 8619, 8621, and 8622.
    • Installed and configured vncserver (TigerVNC) in my LUCI simulator and isola-luci (also updated to the latest CentOS 7) for testing vs. X2Go.
  • This week:
    • Create a test case in my simulated TCS for IT 8573.
    • Work on Observer Scheduler GUI (IT 8571, 8529).
    • Test vncserver on isola-luci.
    • I will be taking half days (possibly full days) off work Friday Apr 29 and Monday May 2, and possibly Thursday Apr 28.

BM

  • Last week:
    • PR for IT 8602, copleyctrl out of range movement error reporting (tested in VM)
    • PR for IT 8553, agw_power device wait for operating status (tested with JCG)
    • PR thAOApp cleanup MsgD connection on exit
    • AO SW meeting - planning for interrupted offset tests
  • Monday LBTI tests with Amali and Juan Carlos:
    • IT 8556: "uninterruptible" aoarbitrator commands worked OK. Tested during DEC?
    • IT 8556: interrupted offsets: collected optical loop data with masterdiagnostic fix.
    • IT 8602: bayside out of range movement error tested, worked OK, ready for DEC checkout
    • AdSecArbitrator command thread safety issue debugging
  • Rest of the week:

SH

  • Tucson network problems seem to be resolved
  • Archive meeting
  • UITS CatNet, backup box
  • Created pipeline.pepsi.lbto.org
  • Created solidworker.tucson.lbto.org
    • Played with SAMBA AD, but ended up adding DNS entries to 10.130.134.4 to point to our old 2008 AD
  • Dan Rapoza's SCADA questions, network questions
  • LBC bsci problems
  • Dropped mountain network complaints from VATT, ASC needs to be accessible to them.

next:
  • Core switches
  • Room 507 upgrades
  • Move shell to hv01 and upgrade to a newer Fedora.
Topic revision: r12 - 26 Apr 2022, DavidEdmeades
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback